OpenStudy (mathwiz007):

Please HELP: Evaluate −[4−2(−1+5)2]. Thanks in Advance.

OpenStudy (mathiscancer):

So first you need to take care of the paranthesis

OpenStudy (mathiscancer):

So -[4-2(4)2] Then distribute the "-"

OpenStudy (mathiscancer):

So right now you have -[4-2*4*2] = -4+16 = 12

OpenStudy (mathiscancer):

So your answer is 12! If I helped click on "best answer" :)